STUDIES OF THE TEMPERATURE DEPENDENCE AND EXCITATION OF THE PRODUCTS OF DIS-SOCIATIVE ELECTRON RECOMBINATION OF N_{2}H^{+} AND HCO^{+} IONS

\underline{V. POTERYA}, J. L. McLAIN, L. M. BABCOCK, and N. G. ADAMS, Department of Chemistry, University of Georgia, Athens, GA 30602, USA.

The dissociative recombination (DR) rate constants of N_{2}H^{+}/N_{2}D^{+} and HCO^{+} ions with electrons have been measured as a function of temperature (100 - 500 K) using the Flowing Afterglow Langmuir Probe (FALP) technique. N_{2}H^{+} and HCO^{+} ions are two of the most abundant ions detected in the interstellar medium. The results cannot be expressed either by the power law T^{−0.5} of the direct DR mechanism or the power law T^{−1.5} of the indirect DR mechanism. In order to further understanding of the results, a study of the vibrational state distributions of the N_{2}(B^{3}Π_{g}) electronically excited products for the N_{2}H^{+} and N_{2}D^{+} recombinations at 100 K has been carried out using an atmospheric pressure monochromator. Comparison has been made with previous results obtained at 300 K ^{a}. The substantial enhancement of the vibrational level (ν′ = 6) detected from the N_{2} B state for N_{2}H^{+} recombination over N_{2}D^{+} recombination is consistent with previous results and can be explained by the influence of a tunneling mechanism of DR. Funding from NASA Grant No. NAG5-8951 is gratefully acknowledged.

aJ. M. Butler, L. M. Babcock, and N. G. Adams, Mol. Physics. 16, 81-90 (1997).